OpenStudy (anonymous):

1. Jeremiah is making a shake by mixing two different protein powders, measured in ounces. The strawberry-flavored powder has 4 grams of protein per ounce. The banana-flavored powder has 3 grams of protein per ounce. He wants the drink to have a total of 6 ounces of powder and contain 22 grams of protein. a. Create a system of linear equations to represent the situation that will determine the exact number of ounces needed for each power. b. Based on the system you wrote, which algebraic method of solving systems of equations will you choose to solve the system? Explain your choice

OpenStudy (anonymous):

c. Solve the system algebraically, showing your work. Then interpret the solution by explaining what your solution represents in the context of the problem.

OpenStudy (anonymous):

olve for y in the first equation and get y=6−x then substitute it in to the second equation and get 4x+3(6−x)=22 and solve that one for x

OpenStudy (anonymous):

aight let me try, s that for A??

OpenStudy (anonymous):

If you solve the first one correctly youll get , y=6 -x. Just solve the last one, 4x+3(6-x)=22

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

how about b?

OpenStudy (anonymous):

Sorry not a good maths helper :/

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

ok i see it now chaser!! ma bad

OpenStudy (anonymous):

Your welcome!!! @The_Chosen